OpenStudy (anonymous):

Parabola help again? Find the vertex, focus, directrix, and focal width of the parabola. x = 4y^2

OpenStudy (anonymous):

This one is backwards.. I figured the vertex was (0,0), but that's all I got.. @campbell_st

OpenStudy (anonymous):

the vertex is (0,0)

OpenStudy (campbell_st):

ok so this is a parabola that looks like  the standard form is \[(y - k)^2 = 4a(x - h)\] (h, k) is the focus and a is the focal length so looking at your question the vertex is at the origin Vertex(0, 0) so re-write it \[y^2 = \frac{1}{4} x\] so to find the focal length solve \[\frac{1}{4} = 4a\] the the directrix will be x = -a focus is (h + a, k) or focus (a, 0) hope it helps